Compute your average velocity in the following two cases: (a) You walk 73.2 m at a speed of 1.22 m/s and then run 73.2 m at a speed of 3.05 m/s along a straight track. (b) You walk for 1.00 min at a speed of 1.22 m/s and then run for 1.00 min at 3.05 m/s along a straight track. (c) Graph x versus t for both cases and indicate how the average velocity is found on the graph.Please help me with the graph

Grade:12th pass

1 Answers

Arun
6 years ago
Well keep in mind that 
distance= rate*time 
so time=distance/rate and rate= distance/time. 
Also Average velocity= distance total moved divided by time that passed (time it took to move that distance) 

With this information we can begin 

a) 
1.22m/s=r 
73.2m=d 
t=? 
we can solve with t=d/r and get 
t =73.2/1.22 and we get t=60sec. 
Doing the same for 
3.05 = r 
73.2 = d 
t =? 
t = 73.2m/3.05m/s= 24secs. 
Using the average velocity information we know we went 
73.2m+73.2m in distance 
60 seconds+24 seconds in time so 
146.4m/84secs.= 1.74m/s 
b) We need to find distance now using the previous formulas we can solve 
d = r*t 
d = 1.22m/s*60sec 
d = 73.2m 

and for the other 
d=3.05m/s*60s 
d=183m 

so same situation with average velocity 
256.2m/120secs= 2.135 m/s